OpenStudy (anonymous):

6 cos(x) -3=0 find all x values of x in interval [0, 2pi]

OpenStudy (anonymous):

We have sin2x=cosx 2sinxcosx−cosx=0 so cosx(2sinx−1)=0 either cos x=0 or 2 sin x-1=0=> sin x =1/2 cos x=0 x=π/2,3π/2 sin x=1/2 x=π/6,(π−π/6)=>π/6,5π/6 so x=π/6,π/2,5π/6,3π/2

OpenStudy (anonymous):

That is incorrect

OpenStudy (tkhunny):

@nelly That is an insufficient response. Though I too wonder why km081 seems to be solving some other problem, you still have shown no work at all. One has to wonder why you know or think the response incorrect. Please prove the answer you already know by solving the problem. 6 cos(x) -3=0 6 cos(x) = 3 cos(x) = 1/2 Go!

OpenStudy (anonymous):

I typed the answers that she provided me and it said it was incorrect